3.Evaluate ∫(3x−2)^6 dx
a.(3x+2)^7
--------- +c
2
b.(3x+2)^7
--------- +c
21

c. (3x-2)^7
--------- +c
21

d.3(3x+2)^7
--------- +c
2

4.Integrate with respect to x :∫3 x
-1 -------dx
√7+x^2
a.2√2
b.4√2
c.4−2√2
d.√2
1
Expert's answer
2019-04-17T04:16:06-0400

Task 3.

"\int(3x-2)^6dx = (1/3)* \int(3x-2)^6d(3x-2)"

"=(1\/3)*(1\/7)*(3x-2)^7 + c,"

c is an integration constant.

Task 4.

"\int3x-1dx=(3/2)*x^2-x+c,"

c is an integration constant.